OpenStudy (anonymous):

How do you find critical points? (get at least 2 of the 3 acceptable answers)

OpenStudy (anonymous):

are you asking a method?

OpenStudy (anonymous):

yeah, for example I know that you can find critical points by taking the derivative and setting it equal to zero, but what are the other possible ways?

OpenStudy (anonymous):

You can plot a graph and then see the critical points - that's another method.

OpenStudy (anonymous):

There's one more way, but I cant seem to get it. :/

OpenStudy (anonymous):

yes..or plot on a sign diagram...(never liked that way) for example, f(x)=(x-1)^2 factor out f(x)=(x=1)*(x-1), set x+1=0 set x-1=0. these will be crit. points

OpenStudy (anonymous):

oops..f(x)=(x+1)*(x-1)

OpenStudy (anonymous):

So, to find critical points are you saying we just factor out the function and set it equal to zero (instead of taking the derivative?)

OpenStudy (anonymous):

my mistake...been a while. no done for first derivative...

OpenStudy (anonymous):

are you trying to get the relative max and min, or also the absolutes?

OpenStudy (anonymous):

For this question?

OpenStudy (anonymous):

no..just trying to get context, since the 2nd der. will give absolutes

OpenStudy (anonymous):

I'm sorry, Im confused.

OpenStudy (anonymous):

well..me too. you want 3 ways, in general, to find the critical values?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

the critical points are max, min and points of inflexion.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok. 2nd derivatives are the inflection points

OpenStudy (anonymous):

1. Find the derivative and set it to 0. This way, you get max and min. The second derivative should give you points of inflexion. 2. Plot the equation on a graph and find the max, min and inflexion points visually. 3. Use a computer/calculator ;)?

OpenStudy (anonymous):

calculator. hehehe

OpenStudy (anonymous):

lol, maybe so. Thanks, guys (:

OpenStudy (anonymous):

good luck